The front of a house is the shape of a triangle. Suppose the base of the triangle is 1m less than its height and the area of the triangle is 15m^2, Find the height of the triangle. (Hint: Use A=1/2bh.)
Answer by mananth(16949)

height =x
base = (x-1)
Area = 1/2 * bases * height
15 = 1/2 * x(x-1)
multiply by 2
30 = x(x-1)
30= x^2-x
x^2-x-30=0
x^2-6x+5x-30=0
x(x-6)+5(x-6)=0
(x-6)(x+5)=0
x=6 OR -5
ignore the negative
height = 6 m
base = 5 m